OpenStudy (anonymous):

x + 1/3 < 4 please solve.

OpenStudy (anonymous):

\( x \lt \frac{11}{3} \)

OpenStudy (anonymous):

thanks. :D

OpenStudy (anonymous):

what is the solution set for this?

OpenStudy (anonymous):

\(\{ x: x \in \mathbb{R} \text{ and } x \in (-\infty, \frac{11}{3}) \} \)